OpenStudy (anonymous):

Simplify: x^2-10x+9/x^2+2x-3

OpenStudy (anonymous):

x^2-10x+9 = x^2 -x -9x + 9 = x(x - 1) -9(x - 1) = (x - 9)(x - 1) x^2+2x-3 = x^2 + 3x - x - 3 = x(x + 3) -1(x + 3) = (x + 3)(x - 1) so x^2-10x+9 (x - 9)(x - 1) x - 9 ----------- = ------------- = ------- x^2+2x-3 (x + 3)(x - 1) x + 3

OpenStudy (anonymous):

tank you

OpenStudy (anonymous):

thank**

OpenStudy (anonymous):

u r welcome...☺